OpenStudy (lgbasallote):

the half-life of carbon-14 is 5,600 years. If 3% of C-14 is found today, find the age of the element. If 2% remains now, find the age of the element.

OpenStudy (anonymous):

\[(\frac{1}{2})^{\frac{t}{5,600}}=.03\] solve for \(t\)

OpenStudy (lgbasallote):

is that t/5600?

OpenStudy (anonymous):

you got this ? it take two steps

OpenStudy (anonymous):

\[\large (\frac{1}{2})^{\frac{t}{5,600}}=.03\]

OpenStudy (lgbasallote):

yeah..i wanted to make sure i didnt do any careless mistakes.

OpenStudy (lgbasallote):

ahh the same as i got! wonderful

OpenStudy (anonymous):

first change of base gives \[\frac{t}{56,000}=\frac{\ln(.03)}{\ln(.5)}\] and then \[t=56,000\times \frac{\ln(.03)}{\ln(.5)}\] and then a calculator

OpenStudy (lgbasallote):

im guessing for the other one it will be \[\huge (\frac 12)^{t/5600} = 0.02?\]

OpenStudy (anonymous):

although these days i guess you can do it calculator in one step,

OpenStudy (anonymous):

yup

OpenStudy (lgbasallote):

ahh wonderful. it seems i was right...i think...i dont recall my answers and yeah we can use calculator for one step now..a lot has changed ever since you entered that cryogenic chamber